40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Sun Oct 10 18:28:25 2010, kiwamu, Summary, SUS, ITMX OSEM offsets OSEM.png
    Reply  Mon Oct 11 10:51:36 2010, steve, Update, SUS, OSEMs, OSEMs, OSEMs...those lovely little OSEMs 40dll.jpg
Message ID: 3686     Entry time: Sun Oct 10 18:28:25 2010     Reply to this: 3688
Author: kiwamu 
Type: Summary 
Category: SUS 
Subject: ITMX OSEM offsets 

Because of the in-vac work on Oct. 4th (see this entry) , ITMX's OSEM offsets were changed.

The two upper OSEMs are still fine, but LL and LR seem to be out of the OSEM's range. 

The plot below shows the trends of LL's and LR's readouts for about two weeks. (The channel name are in the old convention, i.e. ITMY)

OSEM.png

 Some data were missing due to the upgrade of the frame builder.

 It is apparent that the offsets are changed after the in-vac work on Oct. 4th, and now they just show almost zero numbers.

 

The damping of ITMX can still work, if LL and LR are disabled.

At some point before pumping down, we have to check the leveling of the ITMX table again.

ELOG V3.1.3-